OpenStudy (boo.cx):

Point C is the center of the circle. If angle ACB measures (5x + 10)° and arc AB measures (x + 50)° , find x. 7 8 9 10

OpenStudy (babu):

5x + 10 = x + 50 ⇒ x = 10 Hopefully, this should be it